OpenStudy (anonymous):

How do you solve this fraction equation. I think it has something to do with having to flip them but I'm not sure?: t/1 - 1/2 = 1 (Those are not division signs, they are fractions)

OpenStudy (anonymous):

actually they ARE division signs :)

OpenStudy (anonymous):

YES WELL THEY ARE NOT MEANT TO BE.

OpenStudy (anonymous):

t-1/2 = 1 t=3/2

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Fractions are division (numerator/denominator)

OpenStudy (anonymous):

didn't you hear? they aren't supposed to be!

OpenStudy (anonymous):

Well if your into multiplying inverses??

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

good times

OpenStudy (anonymous):

That is what the equation looks like.

OpenStudy (anonymous):

t=9/2

OpenStudy (anonymous):

Thank you. That helped a lot.

OpenStudy (anonymous):

t/3-3/2=0 cross multiply 2t-9=0 solve for t

OpenStudy (anonymous):

Oh okay that makes sense. Thank you. I thought it was something along the lines of that.